Pik-Fast, a chain of convenience stores, is planning to open a new store.  The activities required, their immediate predecessors, and the optimistic, most likely, and pessimistic estimates of their times (in days) are provided below.

Activity Description Immediate Predecessors Optimistic Time (days) Most Likely Time (days) Pessimistic Time (days) TE ES EF LF LS Slack Standard Dev Variance
A Select manager -- 10 12 20 13 0 1.67
B Select site -- 8 15 22 15 0 2.33
C Hire workers A 17 21 31 22 0 2.33
D Basic training C 3 4 5 4 0 0.33
E Secure permits B 12 18 30 19 0 3.00
F Construct building E 26 30 40 31 0 2.33
G Purchase equipment -- 30 45 60 45 0 5.00
H Install equipment F,G 4 6 8 6 0 0.67
I Advertise opening B 15 55 89 54 0 12.33
J Final training D,H 2 3 4 3 0 0.33
K Stock store H,I 4 6 8 6 0 0.67
L Open store J,K 1 1 1 1 0 0.00

b) Identify the critical path and projected project completion time using the expected completion times.

c)  Calculate the expected slack times for any two paths not including the critical path. For the path slack time, compute this compared to the critical path and that also considers the slack time for each activity on that path.

d) Find the probability that the critical path will be completed within 80 working days.

e)  Illustrate or explain why the probability that the critical path will be finished in 80 days or less is not necessarily the probability that the project will be finished in 80 days or less.  Assume that times for each activity will not change from what is provided below and no new activities are added nor are the predecessor relationships changed.
Answer #1

For each activity, calculate the expected time u-(a 4m b)/6 For each activity, Variance o -(b-a)/36. Standard deviation of ea

b)

Critical path is the longest path. It can be identified in the following table. All the activities having Slack time of 0 constitute the critical path.

So, critical path is: B-E-F-H-K-L

Expected completion time = LF time of the last activity of critical path = 78 days

c)

The expected slack times are shown in the right-most last column of the above table

d)

Expected time of critical path = 78 days

Variance of time of critical path = 5.44+9+5.44+0.44+0.44+0 = 20.78

Standard deviation of time of critical path = sqrt(20.78) = 4.56

Expected time of z value = (80-78)/4.56 = 0.4388

F(z) = NORMSDIST(0.4388) = 0.6696

Probability of completion of critical path within 80 working days = 0.6696 ~ 67 %

e) Probability of that the critical path will be finished within 80 days is 0.6696, whereas probability that all the paths will be finished within 80 days is 0.5796. Both are different.

Step 1: List all possible paths through the network alongwith their respective expected times and variances The expected time